OpenStudy (anonymous):

A swimming pool has a volume of 2400 ft^3 and is filled to 8 feet. The owners are draining the pool at 30 ft^3 per minute. After how long is the pool 6 feet deep?

OpenStudy (amistre64):

what assumption should we make about the shape of the pool?

OpenStudy (anonymous):

it doesn't tell us :(

OpenStudy (anonymous):

A) 7.5 minutes Eliminate B) 10 minutes C) 20 minutes D) 30 minutes

OpenStudy (anonymous):

ignore eliminate on A)

OpenStudy (amistre64):

8*bw = 2400 bw = 2400/8 6*bw = volume reminaning how much volume has been removed?

OpenStudy (amistre64):

it seems to me that 600 has been removed at a rate of 30 each minuite how many 30s are in 600?

OpenStudy (anonymous):

I just worked it out. would the answer be C?

OpenStudy (amistre64):

i cant be certain, but yeah .. C is the most likelyone to me

OpenStudy (anonymous):

thank you! i appreciate it! :)

OpenStudy (amistre64):

good luck